# Massively Parallel Minimum Spanning Tree in General Metric Spaces

### v1

@@ Abstract
We study the minimum spanning tree (MST) problem in the massively parallel computation (MPC) model. Our focus is particularly on the strictly sublinear regime of MPC where the space per machine is $O(n^{\delta})$. Here $n$ is the number of vertices and constant $\delta\in(0,1)$ can be made arbitrarily small. The MST problem admits a simple and folklore $O(\log n)$-round algorithm in the MPC model. When the weights can be arbitrary, this matches a conditional lower bound of $\Omega(\log n)$ which follows from a well-known 1vs2-Cycle conjecture. As such, much of the literature focuses on breaking the logarithmic barrier in more structured variants of the problem, such as when the vertices correspond to points in low- [AndoniNikolov, STOC’14] or high-dimensional Euclidean spaces [jayaram2024massively, SODA’24].
In this work, we focus more generally on metric spaces. Namely, all pairwise weights are provided and guaranteed to satisfy the triangle inequality, but are otherwise unconstrained. We show that for any $\varepsilon>0$, a $(1+\varepsilon)$-approximate MST can be found in $O(\log\frac{1}{\varepsilon}+\log\log n)$ rounds, which is the first $o(\log n)$-round algorithm for finding any constant approximation in this setting. Other than being applicable to more general weight functions, our algorithm also slightly improves the $O(\log\log n\cdot\log\log\log n)$ round-complexity of [jayaram2024massively, SODA’24] and significantly improves its approximation from a large constant to $1+\varepsilon$.
@@
It is also worth noting that while many existing lower bounds in the MPC model under the 1vs2-Cycle conjecture only hold against “component stable” algorithms, our lower bound applies to all algorithms. Indeed, a conceptual contribution of our paper is to provide a natural way of lifting the component stability assumption which we hope to have other applications.
@@ 1 Introduction
The minimum spanning tree (MST) problem is one of the most fundamental problems in combinatorial optimization, whose algorithmic study dates back to the work of Borůvka in 1926 [boruuvka1926jistem]. Given a set of points and distances between the points, the goal is to compute a minimum-weight tree over the points. The MST problem has received a tremendous amount of attention from the algorithm design community, and has been studied in a wide variety of computational models, leading to a large toolbox of methods for the problem [Charikar:2002, IT03, indyk2004algorithms, 10.1145/1064092.1064116, har2012approximate, AndoniNikolov, andoni2016sketching, bateni2017affinity, czumaj2009estimating, CzumajEFMNRS05, CzumajS04, chazelle00, chazellerubinfeld].
To deal with the sheer size of these modern embedding datasets, the typical approach is to implement algorithms in massively parallel computation systems such as MapReduce [dean2004mapreduce, dean2008mapreduce], Spark [zaharia2010spark], Hadoop [white2012hadoop], Dryad [isard2007dryad] and others. The Massively Parallel Computation (MPC) model [karloff2010model, goodrich2011sorting, beame2017communication, AndoniNikolov] is a computational model for these systems that balances accurate modeling with theoretical elegance.
@@
The MST problem, in particular, has been extensively studied in the MPC model [AndoniNikolov, bateni2017affinity, karloff2010model, andoni2018parallel, lattanzi2011filtering, Nowicki-STOC21, ahanchi2023massively, chen2022new, jayaram2024massively]. For arbitrary edge-weights, algorithms that obtain an exact MST in $O(\log n)$ rounds are known using connected components algorithms [karloff2010model, andoni2018parallel, behnezhad2019near, coy2022deterministic]. While an $O(\log n)$-round algorithm is often considered too slow in the MPC model (see e.g. [andoni2018parallel, behnezhad2019near] and the references therein), improving the round complexity to sublogarithmic is unlikely. In particular, it is not hard to see that such an algorithm would refute the following widely believed 1vs2-Cycle conjecture [yaroslavtsev2018massively, roughgarden2018shuffles, andoni2018parallel, behnezhad2019near, assadi2019massively]:
@@
[…]
Unfortunately, the progress on the Euclidean version of the problem does not aid in the construction of algorithms for more general metric spaces where the input specifies the pairwise weights $w:\binom{n}{2}\to\mathbb{R}_{+}$ and these weights are only guaranteed to satisfy the triangle inequality. In particular, both the algorithms of [jayaram2024massively] and [AndoniNikolov] crucially rely on the construction of space partitions of Euclidean space, via $\varepsilon$-nets for low-dimensional space, and hyper-grids decompositions for high-dimensional space. These partitions can be efficiently constructed in the MPC model, and are crucial to avoiding $O(\log n)$ round algorithms, as they allow one to bypass running complete connected component algorithms by exploiting the geometry of the space. However, in the general metric case, it is not clear how to exploit the geometry, as we have neither bounded size $\varepsilon$-nets or natural counterparts to hypergrid partitions. Thus it is not clear whether the hardness of the MST problem extends to metric inputs, or whether, as for the Euclidean case, the metric MST problem admits sublogarithmic round algorithms. Specifically, we address the following question:
@@ 1.1 Our Contribution
Other than being applicable to more general metric space, the algorithm of Theorem 1 also improves over the algorithm of [jayaram2024massively] by shaving off a $(\log\log\log n)$ term from its round-complexity and improving its approximation from a large constant to $(1+\varepsilon)$.
@@
When $d\ll n$, our Theorem 1 uses more total space than the $d$-dimensional Euclidean space algorithms of [jayaram2024massively, AndoniNikolov]. However, since the input has size $\Theta(n^{2})$ for general metrics (as opposed to $O(nd)$ for $d$-dimensional Euclidean spaces) this is unavoidable. One may wonder if a more succinct representation of distances, say by taking shortest path distances over a sparse base graph, can reduce total space while keeping the round-complexity sublogarithmic. This is, unfortunately, not possible since the 1vs2-Cycle instance as the base graph implies that $\Omega(\log n)$ rounds are necessary to tell if the MST cost is $12n-$ 1 or $\infty$.
@@
Assuming the 1vs2-Cycle conjecture, Theorem 2 has the following implications:
  • It shows that the $\varepsilon$-dependency of our Theorem 1 is optimal. In particular, for $\varepsilon\leq 1/\log n$, the upper and lower bounds of Theorems 1 and 2 match.
  • It shows that a $(1+o(1))$-approximate MST of a general metric cannot be computed in $O(1)$ rounds (under 1vs2-Cycle). Note that this is unlike the algorithm of [AndoniNikolov] which obtains $(1+o(1))$-approximation of a $d$-dimensional Euclidean space in $O(1)$ rounds whenever $d=o(\log n)$. Therefore, computing MST in general metrics is strictly harder than in low-dimensional Euclidean spaces.
@@
Finally, we believe that our techniques in proving Theorem 2 might be of interest for other MPC lower bounds and particularly getting rid of a common “component stability” assumption. We discuss this in more detail in the following Section 1.2.
@@ 1.2 MPC Lower Bounds: Beyond Component Stability
A powerful approach for proving conditional MPC lower bounds for various graph problems is to combine distributed LOCAL lower bounds with the 1vs2-Cycle conjecture [GhaffariKU19, CzumajDP21a]. On the one hand, an $\Omega(t)$ round distributed LOCAL lower bound for a graph problem implies that vertices need to see $\Omega(t)$-hops away to decide on their output for that problem. On the other hand, the 1vs2-Cycle conjecture essentially implies that to discover vertices that are $\Omega(t)$-hops away, one needs $\Omega(\log t)$ rounds in the sublinear MPC model. Thus, intuitively, an $\Omega(t)$ distributed LOCAL lower bound for a problem $\mathcal{P}$ should imply an $\Omega(\log t)$ round lower bound for $\mathcal{P}$ in MPC under the 1vs2-Cycle conjecture.
[…]
We show in Section A.1 that any distributed local algorithm needs at least $\Omega(1/\varepsilon)$ rounds for computing a $(1+\varepsilon)$-approximate maximum matching in $(1,2)$-metrics (when the communication network is the one induced on the weight one edges). Using the framework of [CzumajDP21a], one can lift this to an MPC lower bound, showing that any component stable MPC algorithm needs $\Omega(\log 1/\varepsilon)$ rounds to find a $(1+\varepsilon)$-approximate MST in $(1,2)$-metrics (see Section A.2). But note that our Theorem 2 applies to all MPC algorithms, not just component stable ones. Instead of component stability, our proof of Theorem 2 heavily relies on the fact that if two vertices $u$ and $v$ are isomorphic to one another, an MPC algorithm should have the same output distribution for both. Note that unlike component stability, this assumption comes without loss of generality as one can randomly relabel the vertex IDs and the algorithm cannot distinguish $u$ from $v$. We hope that our techniques for proving Theorem 2 can pave the way for proving conditional (under 1vs2-Cycle conjecture) MPC lower bounds against all MPC algorithms, not just component-stable ones.
@@ 2 Technical Overview
We present informal overviews of our algorithms and lower bounds of Theorems 1 and 2 in Sections 2.1 and 2.2. Since the techniques in our upper and lower bounds are completely different, these two sections are independent can be read in any order.
2.1 Technical Overview of Theorem 1
@@
We start with a by-now-standard framework for computing MST based on threshold graphs [chazellerubinfeld, CzumajS04, Nowicki-STOC21, jayaram2024massively] and then discuss the challenges that arise in using it in the MPC model.
@@
Given a metric space over $n$-point set $V$ where the distance between $u,v\in V$ is denoted as $w(u,v)\in[1,W]$ ($W=\textnormal{poly}(n)$), the MST of $V$ can be obtained by the following simple variant of the Kruskal’s algorithm. Let $G_{\leq t}$ be a threshold graph over $V$ where there is an edge between vertices $u,v\in V$ iff $w(u,v)\leq t$. Let $\mathcal{P}_{t}$ be the partition of $V$ where each cluster of $\mathcal{P}_{t}$ is a connected component of $G_{\leq t}$. Then $\mathcal{P}_{1},\mathcal{P}_{2},\mathcal{P}_{4},\ldots,\mathcal{P}_{W}$ form a hierarchy of partitions for $V$, i.e., if $u$ and $v$ are in the same cluster of $\mathcal{P}_{t}$, they are also in the same cluster of $\mathcal{P}_{t^{\prime}}$ for all $t^{\prime}\geq t$, and $\mathcal{P}_{W}$ contains only one cluster containing all points of $V$.
@@
Consider each cluster $C$ of $\mathcal{P}_{t}$ for each $t\geq 1$. Suppose $C$ is the union of components $C_{1},C_{2},\ldots,C_{r}$ in $\mathcal{P}_{t/2}$ (define $\mathcal{P}_{1/2}$ as the partition where cluster is a singleton). We regard each $C_{i}$ as a “super node” and compute an MST over $C_{1},C_{2},\ldots,C_{r}$. Then the union of all edges found by the above process for all $t$ and $C$ form a global MST of $V$. Note that every edge between $C_{i}$ and $C_{j}$ in $G_{\leq t}$ must be in the range $[t/2,t]$. This implies that if we compute any arbitrary spanning tree over $C_{1},C_{2},\ldots,C_{r}$ in $G_{\leq t}$, and take the union of all edges found, it gives a $2$-approximate MST. The approximation can be further improved to $(1+\varepsilon)$ by increasing the thresholds by a factor of $(1+\varepsilon)$ instead of $2$.
@@
The problem with the framework above is that under the 1vs2-Cycle conjecture, it is actually impossible to construct the connected components of each threshold graph in $o(\log n)$ rounds exactly (the graph $G_{\leq 1}$ may itself be the 1vs2-Cycle instance). We first overview the approach of [jayaram2024massively] which essentially approximates these components in the Euclidean space via “shifted grids”. We then show how we go beyond shifted grids (which only exist for Euclidean spaces and not general metrics) and solve the problem in general metrics.
MST in Euclidean Spaces via Shifted Grids: An Overview of [jayaram2024massively]
@@
Motivated by the above approximate MST algorithm, [jayaram2024massively] made the following observation: Suppose $\widehat{\mathcal{P}}_{1},\widehat{\mathcal{P}}_{2},\widehat{\mathcal{P}}_{4},\ldots,\widehat{\mathcal{P}}_{W}$ is a randomized hierarchical partitioning of $V$ such that each cluster of $\widehat{\mathcal{P}}_{t}$ is a subset of some connected component of $G_{\leq t}$, and $\forall u,v\in V,t\geq 1$, the probability that $u$ and $v$ are in different clusters of $\widehat{\mathcal{P}}_{t}$ is at most $w(u,v)/(t\cdot\textnormal{poly}(\log n))$. They proved that $\widehat{\mathcal{P}}_{t}$ is a suitable replacement for the set of connected components of $G_{\leq t}$. That is, if we consider each cluster $C$ of $\widehat{\mathcal{P}}_{t}$ for each $t\geq 1$, where $C$ is composed of clusters $C_{1},C_{2},\ldots,C_{r}$ of $\widehat{\mathcal{P}}_{t/2}$, and find an arbitrary spanning tree in $G_{\leq t}$ connecting the super-nodes $C_{1},\ldots,C_{r}$, then the union of all edges found by the above process gives an $O(1)$-approximate MST.
@@
The relaxed partitioning $\widehat{\mathcal{P}}_{t}$ does not increase the MST cost too much because the probability of cutting an MST edge $(u,v)$ at $\widehat{\mathcal{P}}_{t}$ is at most $w(u,v)/(t\cdot\textnormal{poly}(\log n))$, and cutting $(u,v)$ at $\widehat{\mathcal{P}}_{t}$ will introduce an additional cost of at most $t$. Therefore, the overall expected overhead introduced by $(u,v)$ across all levels is at most $\sum_{t=1,2,4,\ldots,W}t\cdot w(u,v)/(t\cdot\textnormal{poly}(\log n))=w(u,v)/\textnormal{poly}(\log n)$. Therefore the problem becomes finding such hierarchical partitioning $\widehat{\mathcal{P}}_{1/2},\widehat{\mathcal{P}}_{1},\widehat{\mathcal{P}}_{2},\ldots,\widehat{\mathcal{P}}_{W}$.
@@
To find such hierarchical partitioning, [jayaram2024massively] proposed the following 2-phase framework:
@@
    Construct $\hat{P}_{t}$ for pivots $t=1/\alpha,1,\alpha,\alpha^{2},\ldots,W$: Here $\alpha$ is a sufficiently large $\textnormal{poly}(\log n)$. Firstly, an auxiliary hierarchical partitioning $\mathcal{P}^{\prime}_{1/\alpha},\mathcal{P}^{\prime}_{2/\alpha},\mathcal{P}^{\prime}_{4/\alpha}\ldots,\mathcal{P}^{\prime}_{W}$ is computed using randomly shifted hypergrids in Euclidean space. The diameter of each cluster of $\mathcal{P}^{\prime}_{t}$ is at most $t$, and the probability that two points $u,v$ are not in the same cluster of $\mathcal{P}^{\prime}_{t}$ is at most $w(u,v)\cdot\textnormal{poly}(\log n)/t$.
    Then, for each $t=1,\alpha,\alpha^{2},\ldots,W$, the partition $\widehat{\mathcal{P}}_{t}$ is obtained by computing the connected components of $G_{\leq t}$ after removing the edges that cross $\mathcal{P}^{\prime}_{\alpha t}$. To compute such connected components, one can start with merging each cluster in $\mathcal{P}^{\prime}_{t}$ as a super-node since each cluster in $\mathcal{P}^{\prime}_{t}$ has diameter at most $t$, and then compute connected components on the edge-removed $G_{\leq t}$. The probability that an edge $(u,v)$ crosses $\widehat{\mathcal{P}}_{t}$ is equal to the probability that $(u,v)$ crosses by $\mathcal{P}^{\prime}_{\alpha t}$ which is at most $w(u,v)\cdot\textnormal{poly}(\log n)/(\alpha t)=w(u,v)/(t\cdot\textnormal{poly}(\log n))$. This property is crucial to the analysis of the approximation ratio.

### v2

@@ Abstract
We study the minimum spanning tree (MST) problem in the massively parallel computation (MPC) model. Our focus is particularly on the strictly sublinear regime of MPC where the space per machine is $O(n^{\delta})$. Here $n$ is the number of vertices and constant $\delta\in(0,1)$ can be made arbitrarily small. The MST problem admits a simple and folklore $O(\log n)$-round algorithm in the MPC model. When the weights can be arbitrary, this matches a conditional lower bound of $\Omega(\log n)$ which follows from a well-known 1vs2-Cycle conjecture. As such, much of the literature focuses on breaking the logarithmic barrier in more structured variants of the problem, such as when the vertices correspond to points in low- [ANOY14, STOC’14] or high-dimensional Euclidean spaces [JMNZ24, SODA’24].
In this work, we focus more generally on metric spaces. Namely, all pairwise weights are provided and guaranteed to satisfy the triangle inequality, but are otherwise unconstrained. We show that for any $\varepsilon>0$, a $(1+\varepsilon)$-approximate MST can be found in $O(\log\frac{1}{\varepsilon}+\log\log n)$ rounds, which is the first $o(\log n)$-round algorithm for finding any constant approximation in this setting. Other than being applicable to more general weight functions, our algorithm also slightly improves the $O(\log\log n\cdot\log\log\log n)$ round-complexity of [JMNZ24, SODA’24] and significantly improves its approximation from a large constant to $1+\varepsilon$.
@@
It is also worth noting that while many existing lower bounds in the MPC model under the 1vs2-Cycle conjecture only hold against “component-stable” algorithms, our lower bound applies to all algorithms. Indeed, a conceptual contribution of our paper is to provide a natural way of lifting the component stability assumption which we hope to have other applications.
@@ 1 Introduction
The minimum spanning tree (MST) problem is one of the most fundamental problems in combinatorial optimization, the algorithmic study of which dates back to the work of Borůvka in 1926 [Bor26]. Given a set of points and distances between the points, the goal is to compute a minimum-weight tree over the points. The MST problem has received a tremendous amount of attention from the algorithm design community, and has been studied in a wide variety of computational models, leading to a large toolbox of methods for the problem [CCF02, IT03, Ind04, FIS05, HIM12, ANOY14, And+16, Bat+17, CS09, Czu+05, CS04, Cha00, CRT05].
To deal with the sheer size of these modern embedding datasets, the typical approach is to implement algorithms in massively parallel computation systems such as MapReduce [DG04, DG08], Spark [ZCFSS10], Hadoop [Whi12], Dryad [IBYBF07] and others. The Massively Parallel Computation (MPC) model [KSV10, GSZ11, BKS17, ANOY14] is a computational model for these systems that balances accurate modeling with theoretical elegance.
@@
The MST problem, in particular, has been extensively studied in the MPC model [ANOY14, Bat+17, KSV10, ASSWZ18, LMSV11, Now21, AAHKZ23, CJLW22, JMNZ24]. For arbitrary edge-weights, algorithms that obtain an exact MST in $O(\log n)$ rounds are known using connected components algorithms [KSV10, ASSWZ18, BDELM19, CC22]. While an $O(\log n)$-round algorithm is often considered too slow in the MPC model (see e.g. [ASSWZ18, BDELM19] and the references therein), improving the round complexity to sublogarithmic is unlikely. In particular, it is not hard to see that such an algorithm would refute the following widely believed 1vs2-Cycle conjecture [YV18, RVW18, ASSWZ18, BDELM19, ASW19]:
@@
A decade ago, [ANOY14] showed that if the vertices correspond to points in a $d$-dimensional Euclidean space for $d=o(\log n)$, then a $(1+o(1))$-approximate MST can be found in $O(1)$ rounds. For high-dimensional Euclidean spaces, where $d=\Omega(\log n)$, [JMNZ24] showed recently that an $O(1)$-approximate MST can be found in $O\big(\log\log n\cdot\log\log\log n\big)$ rounds, breaking the logarithmic-round barrier.
Unfortunately, the progress on the Euclidean version of the problem does not aid in the construction of algorithms for more general metric spaces where the input specifies the pairwise weights $w:\binom{n}{2}\to\mathbb{R}_{+}$ and these weights are only guaranteed to satisfy the triangle inequality. In particular, both the algorithms of [JMNZ24] and [ANOY14] crucially rely on the construction of space partitions of Euclidean space, via $\varepsilon$-nets for low-dimensional space, and hypergrid-decompositions for high-dimensional space. These partitions can be efficiently constructed in the MPC model, and are crucial to avoiding $O(\log n)$ round algorithms, as they allow one to bypass running complete connected component algorithms by exploiting the geometry of the space. However, in the general metric case, it is not clear how to exploit the geometry, as we have neither bounded size $\varepsilon$-nets or natural counterparts to hypergrid partitions. Thus it is not clear whether the hardness of the MST problem extends to metric inputs, or whether, as for the Euclidean case, the metric MST problem admits sublogarithmic round algorithms. Specifically, we address the following question:
@@ 1.1 Our Contribution
Other than being applicable more generally to metric spaces, the algorithm of Theorem˜1 also improves over the algorithm of [JMNZ24] by shaving off a $(\log\log\log n)$ term from its round-complexity and improving its approximation from a large constant to $(1+\varepsilon)$. Also, we note that Theorem˜1, when combined with existing techniques, readily leads to an algorithm for the $(2+\varepsilon)$-approximation of TSP with the same round and space complexity (see Appendix˜B). This is possible due to the hierarchical structure of the MST computed by the algorithm.
@@
When $d\ll n$, our Theorem˜1 uses more total space than the $d$-dimensional Euclidean space algorithms of [JMNZ24, ANOY14]. However, since the input has size $\Theta(n^{2})$ for general metrics (as opposed to $O(nd)$ for $d$-dimensional Euclidean spaces) this is unavoidable. One may wonder if a more succinct representation of distances, say by taking shortest path distances over a sparse base graph, can reduce total space while keeping the round-complexity sublogarithmic. This is, unfortunately, not possible since the 1vs2-Cycle instance as the base graph implies that $\Omega(\log n)$ rounds are necessary to tell if the MST cost is 2n12n-1 or $\infty$.
@@
Assuming the 1vs2-Cycle conjecture, Theorem˜2 has the following implications:
  • It shows that the $\varepsilon$-dependency of our Theorem˜1 is optimal. In particular, for $\varepsilon\leq 1/\log n$, the upper and lower bounds of Theorems˜1 and 2 match.
  • It shows that a $(1+o(1))$-approximate MST of a general metric cannot be computed in $O(1)$ rounds (under 1vs2-Cycle). Note that this is unlike the algorithm of [ANOY14] which obtains $(1+o(1))$-approximation of a $d$-dimensional Euclidean space in $O(1)$ rounds whenever $d=o(\log n)$. Therefore, computing MST in general metrics is strictly harder than in low-dimensional Euclidean spaces.
@@
Finally, we believe that our techniques in proving Theorem˜2 might be of interest for other MPC lower bounds and particularly getting rid of a common “component stability” assumption. We discuss this in more detail in the following Section˜1.2.
@@ 1.2 MPC Lower Bounds: Beyond Component Stability
A powerful approach for proving conditional MPC lower bounds for various graph problems is to combine distributed LOCAL lower bounds with the 1vs2-Cycle conjecture [GKU19, CDP21]. On the one hand, an $\Omega(t)$ round distributed LOCAL lower bound for a graph problem implies that vertices need to see $\Omega(t)$-hops away to decide on their output for that problem. On the other hand, the 1vs2-Cycle conjecture essentially implies that to discover vertices that are $\Omega(t)$-hops away, one needs $\Omega(\log t)$ rounds in the sublinear MPC model. Thus, intuitively, an $\Omega(t)$ distributed LOCAL lower bound for a problem $\mathcal{P}$ should imply an $\Omega(\log t)$ round lower bound for $\mathcal{P}$ in MPC under the 1vs2-Cycle conjecture.
To make the connection above formal, existing methods [GKU19, CDP21] require an additional assumption that the output of the MPC algorithm on one connected component should be independent of its output on other connected components—a property that is referred to as component stability in the literature. While many MPC algorithms are indeed component-stable, there are natural ones that are not—we refer interested readers to the paper of [CDP21] for examples of such algorithms and other limitations of component stability. So, ideally, we would like to get rid of the component stability assumption. Unfortunately, there is still a huge gap between conditional lower bounds against component-stable and non-stable algorithms. We believe that our techniques in proving Theorem˜2 are a step forward towards proving MPC lower bounds without making the component stability assumption.
We show in Section˜A.1 that any distributed local algorithm needs at least $\Omega(1/\varepsilon)$ rounds for computing a $(1+\varepsilon)$-approximate maximum matching in $(1,2)$-metrics (when the communication network is the one induced on the weight-1 edges). Using the framework of [CDP21], one can lift this to an MPC lower bound, showing that any component-stable MPC algorithm needs $\Omega(\log 1/\varepsilon)$ rounds to find a $(1+\varepsilon)$-approximate MST in $(1,2)$-metrics (see Section˜A.2). But note that our Theorem˜2 applies to all MPC algorithms, not just component-stable ones. Instead of component stability, our proof of Theorem˜2 heavily relies on the fact that if two vertices $u$ and $v$ are isomorphic to one another, an MPC algorithm should have the same output distribution for both. Note that unlike component stability, this assumption comes without loss of generality as one can randomly relabel the vertex IDs and the algorithm cannot distinguish $u$ from $v$. We hope that our techniques for proving Theorem˜2 can pave the way for proving conditional (under 1vs2-Cycle conjecture) MPC lower bounds against all MPC algorithms, not just component-stable ones.
@@ 2 Technical Overview
We present informal overviews of our algorithms and lower bounds of Theorems˜1 and 2 in Sections˜2.1 and 2.2. Since the techniques in our upper and lower bounds are completely different, these two sections are independent and can be read in any order.
2.1 Technical Overview of Theorem˜1
@@
We start with a by-now-standard framework for computing MST based on threshold graphs [CRT05, CS04, Now21, JMNZ24] and then discuss the challenges that arise in using it in the MPC model.
@@
Given a metric space over $n$ vertices $V$ where the distance between $u,v\in V$ is denoted as $w(u,v)\in[1,W]$ ($W=\textnormal{poly}(n)$), the MST of $V$ can be obtained by the following simple variant of the Kruskal’s algorithm. Let $G_{\leq t}$ be a threshold graph over $V$ where there is an edge between vertices $u,v\in V$ iff $w(u,v)\leq t$. Let $\mathcal{P}_{t}$ be the partition of $V$ where each vertex set in $\mathcal{P}_{t}$ is a connected component of $G_{\leq t}$. Then $\mathcal{P}_{1},\mathcal{P}_{2},\mathcal{P}_{4},\ldots,\mathcal{P}_{W}$ form a hierarchy of partitions for $V$, i.e., if $u$ and $v$ are in the same set of $\mathcal{P}_{t}$, they are also in the same set of $\mathcal{P}_{t^{\prime}}$ for all $t^{\prime}\geq t$, and $\mathcal{P}_{W}$ contains only one vertex set containing all the vertices.
@@
For each $t\geq 1$, consider a set $C$ of $\mathcal{P}_{t}$. Suppose $C$ is the union of sets $C_{1},C_{2},\ldots,C_{r}$ from $\mathcal{P}_{t/2}$ (define $\mathcal{P}_{1/2}$ as the trivial partition where each set is a singleton). We regard each $C_{i}$ as a super-node and compute an MST over $C_{1},C_{2},\ldots,C_{r}$. Then, the union of the edges found by the above process for all $t$ and $C$ forms an MST for the whole metric. Note that every edge between $C_{i}$ and $C_{j}$ in $G_{\leq t}$ must be in the range $[t/2,t]$. This implies that if we compute any arbitrary spanning tree over $C_{1},C_{2},\ldots,C_{r}$ in $G_{\leq t}$ (instead of an MST), and take the union of all edges found, it gives a $2$-approximate MST. The approximation can be further improved to $(1+\varepsilon)$ by setting the ratio between the thresholds to $(1+\varepsilon)$ instead of $2$.
@@
The problem with the framework above is that under the 1vs2-Cycle conjecture, it is actually impossible to construct the connected components of each threshold graph in $o(\log n)$ rounds exactly (the graph $G_{\leq 1}$ may itself be the 1vs2-Cycle instance). We first overview the approach of [JMNZ24] which essentially approximates these components in the Euclidean space via shifted grids. We then show how we go beyond shifted grids (which only exist for Euclidean spaces and not general metrics) and solve the problem in general metrics.
MST in Euclidean Spaces via Shifted Grids: An Overview of [JMNZ24]
@@
Motivated by the above MST approximation algorithm, [JMNZ24] made the following observation: Suppose $\widehat{\mathcal{P}}_{1},\widehat{\mathcal{P}}_{2},\widehat{\mathcal{P}}_{4},\ldots,\widehat{\mathcal{P}}_{W}$ is a randomized hierarchical partitioning of $V$ such that $(1)$ each vertex set in $\widehat{\mathcal{P}}_{t}$ is a subset of some connected component of $G_{\leq t}$, and $(2)$ for all $u,v\in V,t\geq 1$, the probability that $u$ and $v$ are in different sets of $\widehat{\mathcal{P}}_{t}$ is at most $w(u,v)/(t\cdot\textnormal{poly}(\log n))$. They proved that $\widehat{\mathcal{P}}_{t}$ is a suitable replacement for the set of connected components of $G_{\leq t}$. That is, if we consider each set $C$ of $\widehat{\mathcal{P}}_{t}$ for $t\geq 1$, where $C$ is composed of sets $C_{1},C_{2},\ldots,C_{r}$ of $\widehat{\mathcal{P}}_{t/2}$, and find an arbitrary spanning tree in $G_{\leq t}$ connecting the super-nodes $C_{1},\ldots,C_{r}$, then the union of all edges found by the above process gives an $O(1)$-approximate MST.
@@
The relaxed partitioning $\widehat{\mathcal{P}}_{t}$ does not increase the MST cost too much because the probability of cutting an MST edge $(u,v)$ at $\widehat{\mathcal{P}}_{t}$ is at most $w(u,v)/(t\cdot\textnormal{poly}(\log n))$, and cutting $(u,v)$ at $\widehat{\mathcal{P}}_{t}$ will introduce an additional cost of at most $t$. Therefore, the overall expected overhead introduced by $(u,v)$ across all levels is at most $\sum_{t=1,2,4,\ldots,W}t\cdot w(u,v)/(t\cdot\textnormal{poly}(\log n))=w(u,v)/\textnormal{poly}(\log n)$. Therefore the problem becomes finding such a hierarchy of partitions $\widehat{\mathcal{P}}_{1/2},\widehat{\mathcal{P}}_{1},\widehat{\mathcal{P}}_{2},\ldots,\widehat{\mathcal{P}}_{W}$.
@@
To find the hierarchical partitioning, [JMNZ24] proposed the following 2-phase framework:
@@
    Construct $\hat{P}_{t}$ for pivots $t=1/\alpha,1,\alpha,\alpha^{2},\ldots,W$: Here $\alpha$ is a sufficiently large $\textnormal{poly}(\log n)$. Firstly, an auxiliary hierarchy of partitions $\mathcal{P}^{\prime}_{1/\alpha},\mathcal{P}^{\prime}_{2/\alpha},\mathcal{P}^{\prime}_{4/\alpha}\ldots,\mathcal{P}^{\prime}_{W}$ is computed using randomly shifted hypergrids in Euclidean space. The diameter of each vertex set of $\mathcal{P}^{\prime}_{t}$ is at most $t$, and the probability that two vertices $u,v$ are not in the same set of $\mathcal{P}^{\prime}_{t}$ is at most $w(u,v)\cdot\textnormal{poly}(\log n)/t$.
    Then, for each $t=1,\alpha,\alpha^{2},\ldots,W$, the partition $\widehat{\mathcal{P}}_{t}$ is obtained by computing the connected components of $G_{\leq t}$ after removing the edges that cross $\mathcal{P}^{\prime}_{\alpha t}$. To compute such connected components, one can start with contracting the vertices in each set of $\mathcal{P}^{\prime}_{t}$ as a super-node since each set of $\mathcal{P}^{\prime}_{t}$ has diameter at most $t$, and then compute connected components on the edge-removed $G_{\leq t}$. The probability that an edge $(u,v)$ crosses $\widehat{\mathcal{P}}_{t}$ is equal to the probability that $(u,v)$ crosses by $\mathcal{P}^{\prime}_{\alpha t}$ which is at most $w(u,v)\cdot\textnormal{poly}(\log n)/(\alpha t)=w(u,v)/(t\cdot\textnormal{poly}(\log n))$. This property is crucial to the analysis of the approximation ratio.